graph the system of equations. then solve: x^2 + y^2=10 and xy=3
Answer by checkley77(12844)

x^2 + y^2=10
xy=3 OR x=3/y
(3/y)^2+y^2=10
9/y^2+y^2=10
(9+Y^4)/Y^2=10
9+Y^4=10Y^2
Y^4-10Y^2+9=0
(Y^2-9)(Y^2-1)=0
(Y+3)(Y-3)(Y+1)(Y-1)=0
Y+3=0
Y=-3 ANS.
Y-3=0
Y=3 ANS.
Y+1=0
Y=-1 ANS
Y-1=0
Y=1 ANS.